# Admitting Ignorance Helps the Video Question Answering Models to Answer

### v1

@@ I Introduction
Videoideo Question Answering (VideoQA) has experienced notable progressions, particularly in the realm of deep learning techniques. These contributions can be broadly categorized into two areas: 1) proposing sophisticated model structures to address specific challenges in VideoQA, such as the use of bi-linear attention mechanisms [1, 2] for video-text alignment or the implementation of conditional graph hierarchies for multi-granular understanding of linguistic concepts [3]; 2) pretraining foundation models on large-scale data to enhance generalization abilities, followed by fine-tuning for various downstream tasks [4, 5, 6, 7, 8]. Despite the progress made through novel model structures and large-scale video-text model pretraining, most of these approaches share a common objective in VideoQA: maximizing the correlation between the answer and the video-question pair. However, this goal has limitations, as it can lead to a dilemma where the model must decide whether to trust the video or the question, especially when video-text alignment is suboptimal. In such cases, the model tends to rely solely on either the video or the question for answer prediction, as modeling the correlation between the answer and the video or question alone is more straightforward.
@@ II-A Video Question Answering
Advanced VideoQA methods have been proposed based on deep neural networks to address the problems that exist in this task [21, 2, 3, 22, 23, 24, 25], which is the extension of single-image visual question answering [26, 27, 28, 29, 30, 31, 32]. For example, a dual-LSTM-based approach with both spatial and temporal attention is proposed in [21]. MASN [2] models each object as a graph node and captures the spatial and temporal dependencies of all objects with graph neural networks. HQGA [3] is developed to model the video as a conditional graph hierarchy to align with the multi-granular nature of questions, achieving remarkable results on MSVD and MSRVTT [33]. The atemporal probe (ATP) [22] is presented to degrade the video-language task to image-level understanding, providing a stronger baseline for image-level understanding in the video-language setting than random frames. In summary, these methods focus on adapting various techniques, such as the attention mechanism [34, 35], graph neural networks [2, 36, 37], memory networks [38, 39], and hierarchical structures [40, 3], for improved performance.
@@ III-A Open-Ended VideoQA (OEQA)
OEQA regards VideoQA as a multi-class classification problem, where the answers are considered as classes, and the models are required to choose from a large answer pool given a video and a question. Concretely, given the video $V$, the question $Q$ and the answer pool $\mathcal{A}$, the model aims to predict the conditional distribution of answers $a$, i.e.,
@@ III-C General Model Structure for VideoQA
IV Admitting Ignorance for VideoQA
@@
Global Replacement (Displacement). Specifically, the question in a video-question pair is replaced with a question from other video-question pairs. Note that we avoid replacing the question with general ones, such as “What is the man doing?”, to prevent the displaced question from remaining meaningful for the video. The aim of global replacement is to compel the model to learn coarse correspondences between the video and the question. In other words, if the model successfully acknowledges its lack of knowledge about the globally-intervened input, it indicates that the model has developed the ability to understand both the question and the video in a coarse manner.
@@
IV-A Admitting Ignorance for MCQA
@@
Besides, to prevent the scenario where $\mathcal{A}$ still contains the correct answer to the intervened question, we replace the $N-1$ wrong options with respect to the original question with options randomly from all options in the training set. That is to say, if the question is intervened, the model would choose from the option set defined as $\mathcal{A}^{\prime\prime}=\{b_{1},\cdots,b_{N-1},a_{i^{*}},a_{N+1}\}$, where $\{b_{i}\}_{i=1}^{N-1}$ are randomly-sampled options, $a_{i^{*}}$ is the correct answer to the original question, and $a_{N+1}=\mathrm{``not~{}given\text{''}}$ ”. The model is forced to select $a_{N+1}$. The reason why we keep $a_{i^{*}}$ in $\mathcal{A}^{\prime\prime}$ is that we use it to confuse the model, and force the model to admits its ignorance instead of still choosing $a_{i^{*}}$ when the question is changed.
@@
IV-B Admitting Ignorance for OEQA
@@
IV-C From Admitting Ignorance to Presenting Answers
@@ 14             $\theta=\theta-\gamma\nabla_{\theta}\mathcal{L}$
Furthermore, our prior experiments show that, when using a fixed probability for replacement during training, the model tends to naively admit its ignorance regarding the challenging video-question pairs in the testing set, resulting in poor accuracy. We assume that the reason is a fixed probability leading to underfitting of the original data. In this case, our ignorance-diminishing curriculum learning framework is necessary for better performance. Another strategy involves learning from the original and intervened pairs simultaneously [9, 11], but such an approach requires much more training time and GPU memory. In contrast, our curriculum learning achieves a satisfactory balance between computational cost and performance. The pseudo code for our framework is shown in Algorithm $1$We use SGD as an example. Other optimizers can also be applied..
IV-D Apply to VQA Models
@@ V-A Experiment Settings
V-A2 Training Details.
We follow the suggested fine-tuning settings of InternVideo, including the learning rate, number of epochs, and batch size. Regarding the proposed training framework, several settings and hyper-parameters are crucial to the final performance. Specifically, we set $p(e)$ in Section IV-C to a quadratically-decreasing function as follows,
$p(e)=\frac{p_{r}}{E^{2}}\left(e-E\right)^{2},e\in[1..E],$ (15)
V-B Comparisons to Existing Methods
@@ V-D More Analysis
What is the Impact of the Hyperparameters? We also demonstrate the impact of two crucial hyperparameters in our method: the initial intervention probability ($p_{r}$) and the ratio of displacement to all interventions. We use InternVideo-B for this study. The accuracy on TGIF-Action and MSVD with respect to various values of $p_{r}$ is presented in Fig. 3(a), while keeping the ratio of displacement fixed at 0.5. It has been observed that a larger $p_{r}$ ($p_{r}>0.5$) harms performance, and we assume this is due to the model becoming excessively biased towards the “unknown” answer. Regarding the ratios of displacement, the results are shown in Fig. 3(b), with $p_{r}$ set to 0.3. The results reveal that for TGIF-Action, the highest accuracy is achieved at 0.7, indicating that, for this dataset, displacement plays a more critical role than perturbation. For MSVD, the model performs best when displacement and perturbation are balanced. We fine-tune these two hyperparameters to optimize performance across all datasets.

### v2

@@ I Introduction
Video Question Answering (VideoQA) has experienced notable progressions, particularly in the realm of deep learning techniques. These contributions can be broadly categorized into two areas: 1) proposing sophisticated model structures to address specific challenges in VideoQA, such as the use of bi-linear attention mechanisms [1, 2] for video-text alignment or the implementation of conditional graph hierarchies for multi-granular understanding of linguistic concepts [3]; 2) pretraining foundation models on large-scale data to enhance generalization abilities, followed by fine-tuning for various downstream tasks [4, 5, 6, 7, 8]. Despite the progress made through novel model structures and large-scale video-text model pretraining, most of these approaches share a common objective in VideoQA: maximizing the correlation between the answer and the video-question pair. However, this goal has limitations, as it can lead to a dilemma where the model must decide whether to trust the video or the question, especially when video-text alignment is suboptimal. In such cases, the model tends to rely solely on either the video or the question for answer prediction, as modeling the correlation between the answer and the video or question alone is more straightforward.
@@ II-A Video Question Answering
Advanced VideoQA methods have been proposed based on deep neural networks to address the problems that exist in this task [21, 2, 3, 22, 23, 24, 25], which is the extension of single-image visual question answering [26, 27, 28, 29, 30, 31, 32]. For example, a dual-LSTM-based approach with both spatial and temporal attention is proposed in [21]. MASN [2] models each object as a graph node and captures the spatial and temporal dependencies of all objects with graph neural networks. Besides, [33] presents an approach for video question answering based on temporal structure modeling. An unsupervised encoder-decoder model is used for visual context learning, and a dual-channel ranking loss is proposed for answering questions. HQGA [3] is developed to model the video as a conditional graph hierarchy to align with the multi-granular nature of questions, achieving remarkable results on MSVD and MSRVTT [34]. The atemporal probe (ATP) [22] is presented to degrade the video-language task to image-level understanding, providing a stronger baseline for image-level understanding in the video-language setting than random frames. In summary, these methods focus on adapting various techniques, such as the attention mechanism [35, 36], graph neural networks [2, 37, 38], memory networks [39, 40], and hierarchical structures [41, 3], for improved performance.
@@ III-A Open-Ended VideoQA (OEQA)
OEQA regards VideoQA as a multi-class classification problem [^1] Although open-ended answers can vary in length from a single word to a full sentence, many commonly used OEQA datasets simplify the task by accepting individual words as answers, and our paper adheres to this traditional, simplified approach for modeling OEQA., where the answers are considered as classes, and the models are required to choose from a large answer pool given a video and a question. Concretely, given the video $V$, the question $Q$ and the answer pool $\mathcal{A}$, the model aims to predict the conditional distribution of answers $a$, i.e.,
@@ III-C General Model Structure for VideoQA
IV The Proposed Method
IV-A Causal Perspective for VideoQA
Similar to IGV [9], we have indeed carried out causal analysis in the context of VideoQA. To illustrate the relationships among the key variables involved in VideoQA, we have designed a causal graph, which is presented in Fig. 2. This causal graph helps to visualize the cause-and-effect connections between four important elements: the input video $V$, the input question $Q$, the causal multimodal feature $C$, and the ground-truth answer $A$. Specifically, a detailed breakdown of the causal relationships depicted in the graph is as follows:
  • $V\rightarrow C\leftarrow Q$: The causal multimodal feature $C$ is determined by the combination of the video $V$ and the question $Q$. In other words, $C$ distills the joint semantic information from both the video and the question. This process is crucial as it extracts the relevant multimodal information needed for answering the question accurately.
  • $C\rightarrow A$: The ground-truth answer $A$ should be inferred from the causal multimodal feature $C$. Since $C$ contains the complete information required to answer the question correctly, the model should rely on $C$ to generate the appropriate answer.
  • $Q\dashrightarrow A$: This represents the spurious causality between the input question $Q$ and the answer $A$. Such spurious correlations often stem from language biases that deep models can easily exploit. For instance, in a question like “What is the baby doing?”, the model might simply guess the answer as “crying” without actually analyzing the video content. This might be due to the presence of a strong spurious correlation within the dataset, or perhaps the model faces difficulties in accurately identifying the baby within the video.
Figure 2: The casual graph of VideoQA.
In our work, our main objective is to break this spurious correlation between $Q$ and $A$. We achieve this by making interventions on the questions and compelling the model to admit its ignorance when dealing with certain questions. Instead of just minimizing the empirical risk based on the original question-answering pairs, we introduce additional data. This extra data serves to encourage the model to truly utilize the causal multimodal feature $C$ for making accurate answer predictions. By doing so, we aim to improve the model’s performance and reduce its reliance on spurious correlations, thereby enhancing the overall quality of the VideoQA system.
IV-B Admitting Ignorance for VideoQA
@@
Global Replacement (Displacement). Specifically, the question in a video-question pair is replaced with a question from other video-question pairs. Note that we avoid replacing the question with general ones, such as “What is the man doing?”, to prevent the displaced question from remaining meaningful for the video. More specifically, in practice, we have set a rule to avoid replacing questions that strictly adhere to the template “What does the [SOMEONE] doing?”. By leveraging the template-based nature of our datasets, we can programmatically scan and detect questions in this format, ensuring that they are not subjected to the displacement strategy [^2] The implementation of this exclusion is facilitated by the nature of the datasets we employ. Our datasets are template-based, which means that the structure of each question is highly organized and predictable. This inherent structure allows us to easily identify general questions.. The aim of global replacement is to compel the model to learn coarse correspondences between the video and the question. In other words, if the model successfully acknowledges its lack of knowledge about the globally-intervened input, it indicates that the model has developed the ability to understand both the question and the video in a coarse manner.
@@
We acknowledge that there is a possibility that the intervened questions may not necessarily lead to an “unknown” answer with respect to the video content. However, we would like to clarify that such cases are extremely infrequent. In the datasets we have used, when videos contain multiple individuals, questions typically include an adjective to uniquely identify the person in question. For instance, instead of a simple question like “What is the man doing?”, the question might be “What is the man in red doing?”. In such scenarios, when we perform local replacement by simply changing the subject (e.g., from “man” to “woman” while keeping the adjective “red”), the new question will usually become unrelated to the video content, thus requiring an “unknown” answer. The inherent characteristics of the datasets we employed ensure that the situation almost never occurs.
@@
IV-C Admitting Ignorance for MCQA
@@
Besides, to prevent the scenario where $\mathcal{A}$ still contains the correct answer to the intervened question, we replace the $N-1$ wrong options with respect to the original question with options randomly from all options in the training set. That is to say, if the question is intervened, the model would choose from the option set defined as $\mathcal{A}^{\prime\prime}=\{b_{1},\cdots,b_{N-1},a_{i^{*}},a_{N+1}\}$, where $\{b_{i}\}_{i=1}^{N-1}$ are randomly-sampled options, $a_{i^{*}}$ is the correct answer to the original question, and $a_{N+1}=\mathrm{``not~{}given\text{''}}$ ”. The model is forced to select $a_{N+1}$.
The reason we keep the correct answer of the original question, $a_{i^{*}}$, in the new option set $\mathcal{A}^{\prime\prime}$ is twofold. First, by including $a_{i^{*}}$, we introduce an element of confusion for the model. Since the question has been changed, the model should not be able to rely solely on the presence of the original correct answer to make a decision. Second, we want to force the model to admit its ignorance about the new state of the question. Instead of simply choosing the original correct answer $a_{i^{*}}$ out of habit or due to the presence of the familiar answer in the option set, the model is now required to consider the option $a_{N+1}=\mathrm{``not~{}given"}$. This way, we can more accurately assess the model’s ability to recognize when it doesn’t have the necessary information to make a proper choice and avoid false positives that could occur if the model just selects the original correct answer despite the question being modified. This approach helps us better understand the model’s true understanding and decision-making process when dealing with questions that have been altered, rather than having the model be influenced by the remnants of the original question’s options.
Note that we sample answer options from an extensive pool that encompasses the combined options of all questions in our dataset. This vast options pool contains a rich variety of semantic elements, including nouns, verbs, adjectives, and more, which significantly diversifies the nature of the options. Given this high level of semantic diversity, the likelihood that the randomly sampled options would accurately correspond to an intervened (perturbed) question is extremely low. As the options cover a wide range of concepts and entities, it becomes statistically improbable for them to match the specific requirements of a modified question.
In our proposed framework, the additional “not given” option is specifically intended to denote the situation where “the answer cannot be inferred from the video”. Given that we have engineered the question-video relationships to eliminate valid answers, the model does not need to differentiate between the case of “the answer does not exist in the option” and “the answer cannot be inferred from the video”. In essence, due to our augmentation method, these two seemingly different scenarios converge into a single case that the “not given” option addresses.
@@
IV-D Admitting Ignorance for OEQA
@@
IV-E From Admitting Ignorance to Presenting Answers
@@ 14             $\theta=\theta-\gamma\nabla_{\theta}\mathcal{L}$
Furthermore, our prior experiments show that, when using a fixed probability for replacement during training, the model tends to naively admit its ignorance regarding the challenging video-question pairs in the testing set, resulting in poor accuracy. We assume that the reason is a fixed probability leading to underfitting of the original data. In this case, our ignorance-diminishing curriculum learning framework is necessary for better performance. Another strategy involves learning from the original and intervened pairs simultaneously [9, 11], but such an approach requires much more training time and GPU memory. In contrast, our curriculum learning achieves a satisfactory balance between computational cost and performance. The pseudo code for our framework is shown in Algorithm 1333We use SGD as an example. Other optimizers can also be applied..
IV-F Apply to VQA Models
@@ V-A Experiment Settings
V-A2 Training Details.
We follow the suggested fine-tuning settings of InternVideo, including the learning rate, number of epochs, and batch size. Regarding the proposed training framework, several settings and hyper-parameters are crucial to the final performance. Specifically, we set $p(e)$ in Section IV-E to a quadratically-decreasing function as follows,
$p(e)=\frac{p_{r}}{E^{2}}\left(e-E\right)^{2},e\in[1..E],$ (15)
V-B Comparisons to Existing Methods
Although our proposed methods have achieved improvements across multiple types of datasets, the enhancements on open-ended questions appear to be relatively modest. There are likely two main reasons for this. Firstly, the spurious correlations between questions and answers in the training sets are rather weak. As our methods are designed in part to break these correlations, the limited strength of such correlations reduces the effectiveness of our approach in enhancing performance on open-ended questions. Secondly, the task of modeling “admitting ignorance” in our framework essentially boils down to an open-set classification problem, which remains a highly challenging and unsolved issue in the field. Currently, our approach of simply adding an additional “unknown” class to the answer set is a rather simplistic solution. We are confident that by employing more sophisticated techniques for open-set classification, we can achieve more significant improvements. This will be a key area of focus in our future research endeavors.
@@ V-D More Analysis
What is the Impact of the Hyperparameters? We also demonstrate the impact of two crucial hyperparameters in our method: the initial intervention probability ($p_{r}$) and the ratio of displacement to all interventions. We use InternVideo-B for this study. The accuracy on TGIF-Action and MSVD with respect to various values of $p_{r}$ is presented in Fig. 4(a), while keeping the ratio of displacement fixed at 0.5. It has been observed that a larger $p_{r}$ ($p_{r}>0.5$) harms performance, and we assume this is due to the model becoming excessively biased towards the “unknown” answer. Regarding the ratios of displacement, the results are shown in Fig. 4(b), with $p_{r}$ set to 0.3. The results reveal that for TGIF-Action, the highest accuracy is achieved at 0.7, indicating that, for this dataset, displacement plays a more critical role than perturbation. For MSVD, the model performs best when displacement and perturbation are balanced. When considering only displacement as an intervention, the overall accuracy is slightly lower than when only perturbation is applied. We hypothesize that this is because perturbation compels the model to learn the fine-grained alignment between videos and questions, which is more vital for this task than the coarse-grained alignment that displacement primarily addresses. Furthermore, as depicted in Fig. 3, the proposed framework exhibits remarkable robustness to the inclusion of hyperparameters. Specifically, it shows that for the initial intervention probability and the ratio of displacement/perturbation, the accuracy undergoes only minimal fluctuations. When we assign appropriate values to these two hyperparameters, the accuracy changes are confined to a narrow range, typically between 0.4% and 0.8%. In the context of various datasets, setting the initial intervention probability at 0.3 and establishing the ratio of displacement to perturbation as 1:1 serves as a promising starting point for achieving satisfactory performance.
TABLE V: The comparisons (%) between the models trained with different intervention probabilities (fixed and our dynamic schedule). The improvement with respective to the original model is also provided.
Prob. Multi-Choice Open-Ended
TGIF-A TGIF-T TGIF-FrameQA MSVD-QA
25% 94.1 +1.2 97.4 +0.4 67.7 +0.2 52.1 +1.0
50% 93.3 +0.1 97.4 +0.4 67.5 +0.1 51.2 +0.1
75% 92.2 -0.7 96.1 -0.9 66.2 -1.2 50.4 -0.7
Dynamic 95.3 +2.4 98.4 +1.4 68.2 +0.8 52.7 +1.6
TABLE VI: The comparisons (%) between the models trained with different schedules for intervention probability. The improvement with respective to the original model is also provided.
Sched. Multi-Choice Open-Ended
TGIF-A TGIF-T TGIF-FrameQA MSVD-QA
Linear 94.0 +1.1 97.5 +0.5 67.7 +0.3 51.6 +0.5
Exponential 95.1 +2.2 98.7 +1.7 68.4 +1.0 52.4 +1.3
Quadratic 95.3 +2.4 98.4 +1.4 68.2 +0.8 52.7 +1.6
What is the Impact of Curriculum Learning? We carried out comprehensive experiments to address the lack of a detailed ablation study on intervention probabilities. We conducted two sets of experiments. The first set focused on fixed intervention probabilities, specifically testing values of 25%, 50%, and 75%. The second set explored alternative schedules, including linear and exponential decay. The results of the fixed-probability experiments are reported in Table V, while the findings from the alternative schedule experiments are shown in Table VI. The results indicate that using a fixed intervention probability yields lower accuracy compared to our dynamic strategy. Moreover, a high intervention proportion of 75% negatively impacts the performance. Among the alternative schedules, the linear decay schedule marginally improves the results. The exponential decay schedule achieves accuracy comparable to that of our proposed strategy. These experiments validate the necessity and effectiveness of our current strategy.
TABLE VII: Results on ImageQA datasets of models with and without our training framework.
Model VQA V1 VQA V2
Yes/No Number Other Yes/No Number Other
SAN [70] 78.54 33.46 44.51 68.89 34.55 43.80
SAN+AIQA 79.32 34.20 45.13 69.55 45.44 44.33
MCB [71] 81.62 34.56 52.16 77.91 37.47 51.76
MCB+AIQA 82.61 35.65 53.22 78.99 38.78 52.45
What is the Impact on ImageQA? We conducted supplementary experiments on ImageQA tasks to validate the generalization of our method. We utilized two well-known ImageQA datasets, VQA V1 [72] and VQA V2 [73]. To thoroughly evaluate our training framework, we adapted two established methods, SAN [70] and MCB [71], integrating them with our proposed approach. The experimental results are presented in Table VII. A clear trend emerges from the table: models trained using our framework consistently outperform their original counterparts. This consistent improvement across different datasets and adapted methods strongly attests to the effectiveness of our approach on ImageQA tasks.
What is the Impact of Difference Ways to Admit Ignorance? As proposed in existing works, there are different methods for compelling models to “admit ignorance.” For instance, IGV [9] and TIGV [11] operate by compelling the predicted distribution to be uniform. On the other hand, EIGV [10] uses contrastive learning to enforce that the representation of videos that do not support the answer is distinctly different from the representation of useful video-question pairs. We conducted an evaluation to assess the effectiveness of our proposed method against existing approaches. To carry out this comparison, we integrated our HQGA into the frameworks of IGV, EIGV, and TIGV. These frameworks represent different strategies for prompting models to acknowledge their lack of knowledge. We then evaluated the performance of these integrated models across multiple datasets, with the results summarized in Table VIII. The experimental results reveal several key findings: 1) Our method demonstrates performance that is on par with EIGV and TIGV. This shows that our approach can achieve competitive results in enabling models to admit ignorance, matching the capabilities of these well-regarded existing methods. 2) Across all the evaluated datasets, our method consistently outperforms IGV. This indicates that our approach provides more effective training signals for the model to recognize and admit its lack of knowledge. In addition, a significant advantage of our method lies in its implementation. Unlike the compared methods (IGV, EIGV, and TIGV), which necessitate the addition of extra modules and multiple branches of video feature extraction, our method simply augments the training data. This minimal modification to the model architecture makes our method more efficient and straightforward to implement.
TABLE VIII: The comparisons (%) between the models trained with different ways to admit ignorance.
Method Multi-Choice Open-Ended
TGIF-A TGIF-T TGIF-FrameQA MSVD-QA
IGV [9] 77.5 84.6 57.7 40.5
EIGV [10] 77.9 85.2 58.0 41.4
TIGV [11] 78.1 85.1 58.4 40.6
AIQA (Ours) 78.0 85.3 58.2 41.2
TABLE IX: The comparisons (%) between VideoLLaMA2 (7B) and the finetuned model with our framework. The improvement is also provided.
Model Multi-Choice Open-Ended
MV-Bench [74] MSVD-QA [34]
VideoLLaMA2 [75] 53.4 71.7
VideoLLaMA2+AIQA 53.9 (+0.5) 72.3 (+0.6)
TABLE X: The comparisons (%) between different answer set design. AIQA means that the answer set does not contain the original correct answer. The improvement with respective to the original model is also provided.
Method Multi-Choice
TGIF-Action TGIF-Transition NExT-QA
AIQA 97.1 (+1.9) 98.8 (+1.7) 56.5 (+1.9)
AIQA 96.7 (+1.5) 98.5 (+1.4) 56.3 (+1.7)
What is the Impact on LLM-based Models? We have incorporated the modern LLM-based model VideoLLaMA2 (specifically the 7B version) [75] into our proposed framework. We fine-tuned VideoLLaMA2 using the original configuration and augmented the training samples with our strategy. To provide evidence of the effectiveness of our framework, we have reported the results of both the original VideoLLaMA2 model and the finetuned version on two datasets: MV-Bench [74] (which is used for multi-choice question answering) and MSVD-QA [34] (which is used for open-ended question answering). The results are presented in Table IX. From the reported results, it is evident that the LLM-based video question answering model (VideoLLaMA2) experiences additional performance improvements when using our training framework. This outcome effectively validates our claim that our proposed method is indeed model-agnostic, as it shows that the framework can enhance the performance of different types of models, including modern LLM-based architectures.
What is the Impact of Keeping the Original Correct Options for MCQA? We conduct an experiment to show the effect of the original correct answers, and the results are shown in Table X. As we can see from the comparison, the model trained with our specific design for answer set gains slightly higher accuracy than the naive one for multi-choice video question answering.